# Cycle-Aware ZZ Crosstalk Mitigation on Quantum Hardware

### v1

@@ Abstract
ZZ crosstalk and decoherence hinder superconducting quantum computing. To enhance parallelism in mitigating ZZ crosstalk, we formulate the problem by integrating quantum cycles and two forms of qubit interference. We then propose CAZZO, a Cycle-Aware ZZ crosstalk Optimization algorithm, which uses a timing-based greedy strategy to schedule gates through cycles within quantum circuits. A novel data structure called Time and Distance Dependency Graph is designed to model gate data dependencies and physical distances from quantum topologies for precise scheduling. Additionally, dynamically punching barriers reduces idle time in quantum circuits, further enhancing parallelism. Simulations show a reduction of up to 37.44% in quantum program cycle (14.19% on average) on various NISQ devices with 53 to 127 qubits. Real-device experiments on IBMQ-Brisbane demonstrate significant acceleration in quantum computing while maintaining fidelity.
@@ 1 Introduction
We optimize quantum gate scheduling using a quantum cycle-aware approach. Gate pulses are mapped to cycle intervals. We balance ZZ crosstalk suppression and execution time. This is formalized as the cycle-aware ZZ crosstalk mitigation problem. We propose CAZZO, an algorithm to dynamically optimize gate scheduling between quantum cycles. Our contributions can be summarized as follows.
@@
  • Simulations on IBM and Google devices show that CAZZO improves total program cycle by up to 37.44% (average 14.19%) over state-of-the-art methods. Our method works well for future complex quantum architectures.
  • Real-device experiments on IBMQ-Brisbane confirm that CAZZO maintains fidelity while significantly accelerating computations compared to pulse-based techniques.
The rest of the paper is organized as follows. Section 2 provides a detailed review of quantum computing basics for quantum gate duration and ZZ crosstalk. Section 3 formally describes the cycle-aware ZZ crosstalk mitigation problem. Section 4 describes the key techniques and tools used in CAZZO. Section 5 presents the complete flow of the CAZZO algorithm. Sections 6 and 7 evaluate the algorithm’s performance on both simulated and real devices and illustrate the results of our algorithm respectively. Section 8 compares with previous work relevant to our research. Finally, Section 9 concludes this work and outlines directions for future work.
@@ 3.2 Cycle-Aware ZZ Crosstalk Mitigation
We demonstrate CAZZO’s scheduling advantages through a concrete example on a $3\times 4$ grid topology (Figure 2 (a-b)). Logical qubits $0\sim 11$ are mapped directly to physical qubits for clarity. Table 1 specifies the gate durations in our case.
@@
CAZZO’s Punching Barrier Technique
Our method introduces partial barriers that release qubits immediately after their gates are complete, while maintaining gate dependencies. Figure 2 (d) shows the key innovations of CAZZO:
@@
This strategic barrier removal reduces the total cycle from 12 to 8 (33% improvement) while maintaining equivalent crosstalk suppression. The optimized schedule uses three layers instead of two, demonstrating CAZZO’s ability to improve parallelism through layer fragmentation.
@@ 4 Our Design
We first introduce the key data structure of the CAZZO algorithm: the Time and Distance Dependency Graph (TDDG), along with the concept of punching barriers. The TDDG is designed to accurately capture the spatio-temporal dependencies between quantum gates, providing a well-structured input essential for the algorithm’s execution.
@@ 5 The Algorithm
In this section, we fully discuss how the TDDG data structure and the punching barrier strategy are efficient in the CAZZO algorithm. The methodology is outlined in Figure 5, with the core process highlighted in yellow and explained in detail in this section. The gate scheduling details are demonstrated with the quantum cycles.
@@ 5.1 Preprocessing
In ZZXSched [12], the insertion of identity gates is a key technique used to mitigate ZZ crosstalk. Identity gates are allocated in quantum circuits for two cases: One is to mitigate a parallel-gate set containing only single-qubit gates, and the other is to transfer active-qubit interference into cross-qubit interference for reducing qubit interaction. However, the engagement of the gate hinders the optimal space for compressing the program cycles. To address this, after ZZXSched generates the initial scheduling of gates with the minimum impact of ZZ crosstalk, all identity gates are systematically ignored from the resulting gate set. The output of ZZXSched, now without identity gates, serves as the input of CAZZO. We apply these gate sets to the creation of TDDG.
@@ 5.3 Gate Scheduling and Barrier Punching
After the creation of the TDDG, we can schedule gates through quantum cycles. This section presents the gate scheduling component of the CAZZO algorithm, which optimizes quantum program cycles by consolidating delayed gates into fewer layers and strategically punching barriers within the circuit.
@@ 27 Procedure FindPredecessors(currentLayer)
As a final step, CAZZO reintroduces identity gates into each set of gates to effectively mitigate ZZ crosstalk. The strategy for applying identity gates follows the same principles as ZZXSched. As shown in Figure 6, two identity gates are applied to transfer active-qubit interference to cross-qubit interference. The red edges indicate cross-qubit interference, while the dashed edges represent active-qubit interference.
Figure 7: An example of the CAZZO algorithm extending from the quantum program in Figure 2 (a).
@@ 5.4 An example
To illustrate the CAZZO algorithm, we provide an example in Figure 7 that includes the creation of the TDDG and the subsequent gate scheduling process. This example considers a quantum circuit with five gates — $\mathrm{iSWAP_{0}}$, $\mathrm{CZ_{0}}$, $\mathrm{RZ_{0}}$, $\mathrm{CZ_{1}}$, and $\mathrm{iSWAP_{1}}$ — executed on nine qubits ($\mathrm{q_{2}},\mathrm{q_{3}},\mathrm{q_{4}},\mathrm{q_{5}},\mathrm{q_{6}},\mathrm{q_{7}},\mathrm{q_{8}},\mathrm{q_{9}},\mathrm{q_{10}}$, as shown in Figure 2) with latencies of 6, 2, and 1 time units. Each gate is represented as a node in the TDDG, with dependencies as directed edges, such as $\mathrm{iSWAP_{0}}\xrightarrow{}\mathrm{CZ_{1}}$ and $\mathrm{iSWAP_{0}}\xrightarrow{}\mathrm{iSWAP_{1}}$, indicating that both gates must follow $\mathrm{iSWAP_{0}}$. Initially, gates are assigned to layers based on dependencies. Layer 1 contains $\mathrm{iSWAP_{0}}$, $\mathrm{CZ_{0}}$, and $\mathrm{RZ_{0}}$, which can be executed in parallel, while Layer 2 contains $\mathrm{CZ_{1}}$ and $\mathrm{iSWAP_{1}}$. The LMFT for the first layer is $\tau_{layer}=6$, and the GEST for $\mathrm{CZ_{1}}$ and $\mathrm{iSWAP_{1}}$ in the following layer is $\tau_{gate}=6$ and $\tau_{gate}=2$, respectively, based on their predecessors’ finish times. The cross-layer gate is $\mathrm{iSWAP_{0}}$, as $\mathrm{iSWAP_{1}}$ is advanceable. Both SWAP gates are placed in PEZ, updating the finish time of the second SWAP gate to 8. In the next iteration, PEZ = $\mathrm{iSWAP_{0}}$, $\mathrm{iSWAP_{1}}$} will be used to determine the next cross-layer gates. The next identified cross-layer gate is $\mathrm{iSWAP_{1}}$.
@@ 6 Evaluation
We evaluated CAZZO using benchmarks on the latest quantum hardware, focusing on both simulations and real-device experiments. The evaluation compares CAZZO’s effectiveness with existing ZZ crosstalk mitigation methods, using execution time reduction and fidelity as key metrics. In this work, our aim is to address the following research questions.
@@
  • $[RQ4]$: How does CAZZO scale with increasing quantum circuit size?
@@
To ensure a fair evaluation of CAZZO, we compare it with standard quantum execution and ZZXSched, an advanced framework for co-optimizing gate scheduling and pulse control to mitigate ZZ crosstalk [12]. ZZXSched includes gate scheduling at the software level and pulse optimization at the hardware level. However, to ensure a fair comparison that isolates the software-based scheduling improvements introduced by CAZZO, we exclude ZZXSched’s pulse optimization component in our experiments. This enables us to directly evaluate the improvements in execution time and parallelism derived purely from CAZZO’s scheduling optimizations.
@@
To evaluate CAZZO’s performance, we use 72 benchmarks from the QASMBench suite [16], a widely recognized collection designed to assess NISQ devices. The benchmarks cover various quantum algorithms, categorized by size: small-scale (2–10 qubits, 11–1008 gates, depths of 2–551), medium-scale (11–27 qubits, 22–2016 gates, depths of 10–2987), and large-scale (28–76 qubits, 40–959 gates, depths of 32–9265).
@@
We implemented our CAZZO scheduling algorithm using Python 3.9, interfacing with the IBM Qiskit software library [17]. The Qiskit transpiler was utilized to compile the logical quantum circuits from the QASMBench benchmarks into executable forms on actual quantum hardware. To ensure that CAZZO’s contribution to scheduling optimization is isolated, we set the Qiskit optimization level to 0, disabling any other transpiler-level optimizations that could interfere with the results of our scheduling algorithm. The compilation process primarily uses the SABRE mapping algorithm [18], which is widely used to map logical qubits to physical ones on a quantum processor. For evaluating performance across different quantum hardware platforms, we adapted the coupling maps and physical basis gates to suit each machine’s architecture.
@@
2) Settings for Real-Device Experiments: We evaluated CAZZO on IBMQ-Brisbane, a 127-qubit superconducting quantum processor for real condition tests. The topology of IBMQ-Brisbane offers a balanced qubit connectivity that helps reduce crosstalk while maintaining high coherence times [21]. Its architecture makes it ideal for testing large-scale quantum circuits.
@@ 6.1 Evaluation Metrics
We evaluated CAZZO by two key metrics: speedup ratio and fidelity. The calculation method is detailed as follows.
@@
To measure the performance gains achie-ved by CAZZO, we calculate the speedup ratio as follows:
$\Delta=\frac{\tau_{ZZXSched}-\tau_{CAZZO}}{\tau_{ZZXSched}}$ (4)
where $\Delta$ represents the speedup ratio, and $\tau_{CAZZO}(\tau_{ZZXSched})$ denotes the total execution time of the quantum circuit on a quantum machine using the CAZZO (ZZXSched) algorithm. A higher $\Delta$ indicates a greater efficiency gain.
$\tau_{CAZZO(ZZXSched)}=\sum_{n=0}^{layers}{t_{layercycle}}$ (5)
Here $\tau_{CAZZO(ZZXSched)}$ is calculated by adding the duration of all layers. By comparing $\tau_{CAZZO}$ and $\tau_{ZZXSched}$, we quantify the improvement in efficiency derived from the CAZZO algorithm’s optimization of gate scheduling and parallelism [14].
@@ 7 Results
This section answers the research questions proposed in Section 6 and presents a detailed analysis of the experimental results obtained from both simulated and real-device tests of CAZZO across various quantum hardware platforms.
@@ 7.1 RQ1 — Circuit Execution Speedup
The primary goal of the CAZZO algorithm is to reduce the execution time of quantum circuits. Our experiments demonstrate that CAZZO substantially outperforms the current state-of-the-art algorithm, ZZXSched, across multiple quantum topologies including IBM, Google, and Rigetti systems.
@@ 7.1.1 IBMQ Devices
Figure 9 (a) illustrates the speedup achieved by CAZZO on IBM’s 127-qubit Brisbane processor. On average, CAZZO reduced the circuit execution time by 14.19% across the set of benchmarks, with the $dnn\_n16$ benchmark showing the most significant improvement, achieving a speedup of 35.85%. This result highlights the efficiency of CAZZO in handling larger circuits, where qubit interactions and gate scheduling become more complex.
@@ 7.1.2 Google Sycamore Device
On Google’s 53-qubit Sycamore processor (Figure 9 (b)), CAZZO achieved an average execution time reduction of 6.02%. Although the speedup on Sycamore was less pronounced than on IBMQ devices, the CAZZO’s performance is still notable, with a maximum speedup of 22.02% observed in certain benchmarks. This difference in performance can be partly attributed to Sycamore’s native gate set, which includes faster gate operations such as the $\sqrt{\mathrm{iSWAP}}$ and CZ gates. As a result, the relative gains from optimizing gate scheduling are smaller compared to devices where gate latencies are more varied.
@@ 7.1.3 Rigetti Aspen-M and Ankaa-Q3 Devices
Rigetti’s Aspen-M and Ankaa-Q3 devices also showed positive results with CAZZO (Figures 9 (c) and 9 (d)). CAZZO achieved an average speedup of 6.27% on Aspen-M and 4.25% on Ankaa-Q3, with the best case on Aspen-M reaching a remarkable 37.44% improvement in execution time. This shows that the CAZZO algorithm is effective in various quantum topologies.

### v2

@@ Abstract
ZZ crosstalk and decoherence hinder superconducting quantum computing. To enhance parallelism in mitigating ZZ crosstalk, we formulate the problem by integrating quantum cycles and two forms of qubit interference. We then propose CYCO, a CYcle-aware ZZ Crosstalk Optimization algorithm, which uses a timing-based greedy strategy to schedule gates through cycles within quantum circuits. A novel data structure called Time and Distance Dependency Graph is designed to model gate data dependencies and physical distances from quantum topologies for precise scheduling. Additionally, dynamically punching barriers reduces idle time in quantum circuits, further enhancing parallelism. Simulations show a reduction of up to 37.44% in quantum program cycle (14.19% on average) on various NISQ devices with 53 to 127 qubits. Real-device experiments on IBMQ-Brisbane demonstrate significant acceleration in quantum computing while maintaining fidelity.
@@ 1 Introduction
We optimize quantum gate scheduling using a quantum cycle-aware approach. Gate pulses are mapped to cycle intervals. We balance ZZ crosstalk suppression and execution time. This is formalized as the cycle-aware ZZ crosstalk mitigation problem. We propose CYCO, an algorithm to dynamically optimize gate scheduling between quantum cycles. Our contributions can be summarized as follows.
@@
  • Simulations on IBM and Google devices show that CYCO improves total program cycle by up to 37.44% (average 14.19%) over state-of-the-art methods. Our method works well for future complex quantum architectures.
  • Real-device experiments on IBMQ-Brisbane confirm that CYCO maintains fidelity while significantly accelerating computations compared to pulse-based techniques.
The rest of the paper is organized as follows. Section 2 provides a detailed review of quantum computing basics for quantum gate duration and ZZ crosstalk. Section 3 formally describes the cycle-aware ZZ crosstalk mitigation problem. Section 4 describes the key techniques and tools used in CYCO. Section 5 presents the complete flow of the CYCO algorithm. Sections 6 and 7 evaluate the algorithm’s performance on both simulated and real devices and illustrate the results of our algorithm respectively. Section 8 compares with previous work relevant to our research. Finally, Section 9 concludes this work and outlines directions for future work.
@@ 3.2 Cycle-Aware ZZ Crosstalk Mitigation
We demonstrate CYCO’s scheduling advantages through a concrete example on a $3\times 4$ grid topology (Figure 2 (a-b)). Logical qubits $0\sim 11$ are mapped directly to physical qubits for clarity. Table 1 specifies the gate durations in our case.
@@
CYCO’s Punching Barrier Technique
Our method introduces partial barriers that release qubits immediately after their gates are complete, while maintaining gate dependencies. Figure 2 (d) shows the key innovations of CYCO:
@@
This strategic barrier removal reduces the total cycle from 12 to 8 (33% improvement) while maintaining equivalent crosstalk suppression. The optimized schedule uses three layers instead of two, demonstrating CYCO’s ability to improve parallelism through layer fragmentation.
@@ 4 Our Design
We first introduce the key data structure of the CYCO algorithm: the Time and Distance Dependency Graph (TDDG), along with the concept of punching barriers. The TDDG is designed to accurately capture the spatio-temporal dependencies between quantum gates, providing a well-structured input essential for the algorithm’s execution.
@@ 5 The Algorithm
In this section, we fully discuss how the TDDG data structure and the punching barrier strategy are efficient in the CYCO algorithm. The methodology is outlined in Figure 5, with the core process highlighted in yellow and explained in detail in this section. The gate scheduling details are demonstrated with the quantum cycles.
@@ 5.1 Preprocessing
In ZZXSched [12], the insertion of identity gates is a key technique used to mitigate ZZ crosstalk. Identity gates are allocated in quantum circuits for two cases: One is to mitigate a parallel-gate set containing only single-qubit gates, and the other is to transfer active-qubit interference into cross-qubit interference for reducing qubit interaction. However, the engagement of the gate hinders the optimal space for compressing the program cycles. To address this, after ZZXSched generates the initial scheduling of gates with the minimum impact of ZZ crosstalk, all identity gates are systematically ignored from the resulting gate set. The output of ZZXSched, now without identity gates, serves as the input of CYCO. We apply these gate sets to the creation of TDDG.
@@ 5.3 Gate Scheduling and Barrier Punching
After the creation of the TDDG, we can schedule gates through quantum cycles. This section presents the gate scheduling component of the CYCO algorithm, which optimizes quantum program cycles by consolidating delayed gates into fewer layers and strategically punching barriers within the circuit.
@@ 27 Procedure FindPredecessors(currentLayer)
As a final step, CYCO reintroduces identity gates into each set of gates to effectively mitigate ZZ crosstalk. The strategy for applying identity gates follows the same principles as ZZXSched. As shown in Figure 6, two identity gates are applied to transfer active-qubit interference to cross-qubit interference. The red edges indicate cross-qubit interference, while the dashed edges represent active-qubit interference.
Figure 7: An example of the CYCO algorithm extending from the quantum program in Figure 2 (a).
@@ 5.4 An example
To illustrate the CYCO algorithm, we provide an example in Figure 7 that includes the creation of the TDDG and the subsequent gate scheduling process. This example considers a quantum circuit with five gates — $\mathrm{iSWAP_{0}}$, $\mathrm{CZ_{0}}$, $\mathrm{RZ_{0}}$, $\mathrm{CZ_{1}}$, and $\mathrm{iSWAP_{1}}$ — executed on nine qubits ($\mathrm{q_{2}},\mathrm{q_{3}},\mathrm{q_{4}},\mathrm{q_{5}},\mathrm{q_{6}},\mathrm{q_{7}},\mathrm{q_{8}},\mathrm{q_{9}},\mathrm{q_{10}}$, as shown in Figure 2) with latencies of 6, 2, and 1 time units. Each gate is represented as a node in the TDDG, with dependencies as directed edges, such as $\mathrm{iSWAP_{0}}\xrightarrow{}\mathrm{CZ_{1}}$ and $\mathrm{iSWAP_{0}}\xrightarrow{}\mathrm{iSWAP_{1}}$, indicating that both gates must follow $\mathrm{iSWAP_{0}}$. Initially, gates are assigned to layers based on dependencies. Layer 1 contains $\mathrm{iSWAP_{0}}$, $\mathrm{CZ_{0}}$, and $\mathrm{RZ_{0}}$, which can be executed in parallel, while Layer 2 contains $\mathrm{CZ_{1}}$ and $\mathrm{iSWAP_{1}}$. The LMFT for the first layer is $\tau_{layer}=6$, and the GEST for $\mathrm{CZ_{1}}$ and $\mathrm{iSWAP_{1}}$ in the following layer is $\tau_{gate}=6$ and $\tau_{gate}=2$, respectively, based on their predecessors’ finish times. The cross-layer gate is $\mathrm{iSWAP_{0}}$, as $\mathrm{iSWAP_{1}}$ is advanceable. Both SWAP gates are placed in PEZ, updating the finish time of the second SWAP gate to 8. In the next iteration, PEZ = $\mathrm{iSWAP_{0}}$, $\mathrm{iSWAP_{1}}$} will be used to determine the next cross-layer gates. The next identified cross-layer gate is $\mathrm{iSWAP_{1}}$.
@@ 6 Evaluation
We evaluated CYCO using benchmarks on the latest quantum hardware, focusing on both simulations and real-device experiments. The evaluation compares CYCO’s effectiveness with existing ZZ crosstalk mitigation methods, using execution time reduction and fidelity as key metrics. In this work, our aim is to address the following research questions.
@@
  • $[RQ4]$: How does CYCO scale with increasing quantum circuit size?
@@
To ensure a fair evaluation of CYCO, we compare it with standard quantum execution and ZZXSched, an advanced framework for co-optimizing gate scheduling and pulse control to mitigate ZZ crosstalk [12]. ZZXSched includes gate scheduling at the software level and pulse optimization at the hardware level. However, to ensure a fair comparison that isolates the software-based scheduling improvements introduced by CYCO, we exclude ZZXSched’s pulse optimization component in our experiments. This enables us to directly evaluate the improvements in execution time and parallelism derived purely from CYCO’s scheduling optimizations.
@@
To evaluate CYCO’s performance, we use 72 benchmarks from the QASMBench suite [16], a widely recognized collection designed to assess NISQ devices. The benchmarks cover various quantum algorithms, categorized by size: small-scale (2–10 qubits, 11–1008 gates, depths of 2–551), medium-scale (11–27 qubits, 22–2016 gates, depths of 10–2987), and large-scale (28–76 qubits, 40–959 gates, depths of 32–9265).
@@
We implemented our CYCO scheduling algorithm using Python 3.9, interfacing with the IBM Qiskit software library [17]. The Qiskit transpiler was utilized to compile the logical quantum circuits from the QASMBench benchmarks into executable forms on actual quantum hardware. To ensure that CYCO’s contribution to scheduling optimization is isolated, we set the Qiskit optimization level to 0, disabling any other transpiler-level optimizations that could interfere with the results of our scheduling algorithm. The compilation process primarily uses the SABRE mapping algorithm [18], which is widely used to map logical qubits to physical ones on a quantum processor. For evaluating performance across different quantum hardware platforms, we adapted the coupling maps and physical basis gates to suit each machine’s architecture.
@@
2) Settings for Real-Device Experiments: We evaluated CYCO on IBMQ-Brisbane, a 127-qubit superconducting quantum processor for real condition tests. The topology of IBMQ-Brisbane offers a balanced qubit connectivity that helps reduce crosstalk while maintaining high coherence times [21]. Its architecture makes it ideal for testing large-scale quantum circuits.
@@ 6.1 Evaluation Metrics
We evaluated CYCO by two key metrics: speedup ratio and fidelity. The calculation method is detailed as follows.
@@
To measure the performance gains achie-ved by CYCO, we calculate the speedup ratio as follows:
$\Delta=\frac{\tau_{ZZXSched}-\tau_{CYCO}}{\tau_{ZZXSched}}$ (4)
where $\Delta$ represents the speedup ratio, and $\tau_{CYCO}(\tau_{ZZXSched})$ denotes the total execution time of the quantum circuit on a quantum machine using the CYCO (ZZXSched) algorithm. A higher $\Delta$ indicates a greater efficiency gain.
$\tau_{CYCO(ZZXSched)}=\sum_{n=0}^{layers}{t_{layercycle}}$ (5)
Here $\tau_{CYCO(ZZXSched)}$ is calculated by adding the duration of all layers. By comparing $\tau_{CYCO}$ and $\tau_{ZZXSched}$, we quantify the improvement in efficiency derived from the CYCO algorithm’s optimization of gate scheduling and parallelism [14].
@@ 7 Results
This section answers the research questions proposed in Section 6 and presents a detailed analysis of the experimental results obtained from both simulated and real-device tests of CYCO across various quantum hardware platforms.
@@ 7.1 RQ1 — Circuit Execution Speedup
The primary goal of the CYCO algorithm is to reduce the execution time of quantum circuits. Our experiments demonstrate that CYCO substantially outperforms the current state-of-the-art algorithm, ZZXSched, across multiple quantum topologies including IBM, Google, and Rigetti systems.
@@ 7.1.1 IBMQ Devices
Figure 9 (a) illustrates the speedup achieved by CYCO on IBM’s 127-qubit Brisbane processor. On average, CYCO reduced the circuit execution time by 14.19% across the set of benchmarks, with the $dnn\_n16$ benchmark showing the most significant improvement, achieving a speedup of 35.85%. This result highlights the efficiency of CYCO in handling larger circuits, where qubit interactions and gate scheduling become more complex.
@@ 7.1.2 Google Sycamore Device
On Google’s 53-qubit Sycamore processor (Figure 9 (b)), CYCO achieved an average execution time reduction of 6.02%. Although the speedup on Sycamore was less pronounced than on IBMQ devices, the CYCO’s performance is still notable, with a maximum speedup of 22.02% observed in certain benchmarks. This difference in performance can be partly attributed to Sycamore’s native gate set, which includes faster gate operations such as the $\sqrt{\mathrm{iSWAP}}$ and CZ gates. As a result, the relative gains from optimizing gate scheduling are smaller compared to devices where gate latencies are more varied.
@@ 7.1.3 Rigetti Aspen-M and Ankaa-Q3 Devices
Rigetti’s Aspen-M and Ankaa-Q3 devices also showed positive results with CYCO (Figures 9 (c) and 9 (d)). CYCO achieved an average speedup of 6.27% on Aspen-M and 4.25% on Ankaa-Q3, with the best case on Aspen-M reaching a remarkable 37.44% improvement in execution time. This shows that the CYCO algorithm is effective in various quantum topologies.